(x-11) squared =45
Found 2 solutions by FireFox, richwmiller:
Answer by FireFox(3)

x=2036
(x-11) squared = 45
= (x-11) = 45^2
= (x-11) = 2025
= x = 2025 + 11
= x = 2036

Answer by richwmiller(17219)

(x-11)^2=45
x-11=+-sqrt(45)
x=11+sqrt(45)
x=11-sqrt(45)
sqrt(45)=sqrt(9*5)=3sqrt(5)
x=11+3*sqrt(5)
x=11-3*sqrt(5)